A rocket climbs vertically and is powered in such a way that it has constant acceleration a. It reaches a height of 1km with a velocity of 100m/s. What is the value of a? How long does the rocket take to reach this 1km height?
1
Expert's answer
2019-07-16T09:11:59-0400

Define height of rocket H = 1 km, velocity v = 100 m/s, time to reach 1 km t.

Since rocket has constant acceleration


"H = v*t\/2,"

"t=2*H\/v=2*1000m\/(100m\/s)=20s"

"a=v\/t=100m\/s\/(20s)=5m\/s^2"

Answer: a=5m/s2, t=20s.